# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ VII.2 Comparison to measurement-based quantum computation
The cluster state—the canonical resource for performing measurement-based quantum computation (MBQC) [7]—is a specific example of a state that can be prepared in constant depth and for which we can guarantee long-ranged MIE (provided we measure in the appropriate basis): By construction, this state has the property that upon measuring qubits in bases within the $x$-$y$ plane, a random unitary circuit is executed in the sideways direction. The specific unitary implemented in this sideways circuit depends on the measurement outcomes $s_{B}$, and thus the post-measurement state received by $A$ and $C$ after this process is given by |$\ket{\phi^{AC}_{s_{B}}}=(\mathbbm{I}\otimes U^{\rm MBQC}_{s_{B}})\ket{\Phi^{+}_{AC}}$, where $U^{\rm MBQC}_{s_{B}}$ describes the unitary circuit conditioned on the outcomes, and |$\ket{\Phi^{+}_{AC}}$ is some maximally entangled state. We can relate the unitarity of this sideways circuit to the fact that such states have maximal measurement-induced entanglement (this is a consequence of the Schrödinger-HJW theorem [95]).
@@
We remark that the picture described here very naturally applies to the problem of measuring cluster states in bases out of the plane, where entanglement transitions have been observed numerically [16, 81]. Additionally, the sideways-evolving circuit picture could be used to analyse strategies for distributing entanglement in quantum networks [103, 104], where similar kinds of short- to long-ranged entanglement transitions can occur.
@@ VII.3 Connections to hardness of sampling
As mentioned in the introduction, there is a close relationship between measurement-induced entanglement in shallow 2D circuits and the hardness of simulating the output distribution of such circuits. Indeed, the known worst-case examples of finite-depth states that are provably hard to sample from [2, 105] (assuming certain widely-believed conjectures in complexity theory) use resource states for MBQC, which as we saw above can exhibit maximal MIE. Here we discuss how our results can be placed within the context of quantum advantage in shallow-depth circuits.
@@ VIII Conclusion and outlook
Our results raise several important open questions. One natural direction would be to consider the effects of experimentally relevant perturbations, such as noise and measurement readout errors. While a constant rate of noise without any change of circuit architecture is likely to lead to result in a distribution that is easy to sample from classically (as is the case in deep random circuit sampling [111]), hardness results for noisy circuit sampling in the constant-depth regime have been put forward in three-dimensional systems [112, 113], and also in certain adaptive settings [114]. One can also ask the separate question of whether post-measurement states in these settings with noise remain long-ranged entangled, which could be addressed by generalizing our distillation-based arguments to mixed states, e.g. using a decoupling approach [64].
The states we have considered in this work are by definition short-ranged entangled. Using our isoTNS representation of these states, we showed how the sampling problem is equivalent to a non-unitary monitored quantum dynamics evolving in the spatial direction. More generally, isometric tensor networks can be used to represent a more general class of wavefunctions, including those with topological order [45]. It would be interesting to understand how the structure of such long-ranged entangled states imprints itself on the effective monitored dynamics, and in turn, how the post-measurement states are themselves affected.
@@ Acknowledgements.
MM acknowledges support from Trinity College, Cambridge. WWH is supported by the Singapore NRF Fellowship, NRF-NRFF15-2023-0008, and the CQT Bridging Fund. DM acknowledges support from the Novo Nordisk Fonden under grant number NNF22OC0071934 and NNF20OC0059939. Discussions between MM and WWH on the subject of this work were initiated in part at the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-1607611.
@@ Appendix A Isometric tensor network representation of finite-depth states
Figure 10: (a) After grouping legs together appropriately, the collections of gates depicted in Fig. 9 can be connected up into the network shown. Black dots representing physical (blocked) qudits, which are arranged into cells, demarcated with red dashed lines. The intra-cell unitaries are left implicit here. (b) If we consider all qudits within a cell as one collective degree of freedom, we can write |$\ket{\Psi_{\rm 2D}}$ as a tensor network state, where each five-legged tensor (c, blue squares) is an isometry from two virtual legs to the physical leg (out of the page) plus the other two virtual legs. This defines as isoTNS state [46], with an orthogonality center in a particular corner of the lattice.
@@ Appendix B Ising partition function bound for holographic model
where [$\mathcal{H}[\{\sigma_{i}\}]$ is equal to the 2-Rényi entropy of the corresponding region $I$ in the pre-measurement state, as we saw in Eq. (22). In the holographic circuit model, $H[\{\sigma_{i}\}]$ is equal to the entropy of the bond state $S_{e}^{(2)}$ times the total number of bonds on which the Ising spins anti-align. The right hand side of (71) then becomes the partition function for the Ising model with opposite boundary conditions on $A$ and $C$, where the inverse temperature is $\beta=S_{e}^{(2)}/2$. We can compare this to $\mathcal{Z}_{+-}$ [Eq. (21)], which we encountered in the calculation of quasientropy: the difference is just a factor of 2 in the effective temperature. By using Cauchy-Schwartz ($(\sum_{\{\sigma_{i}\}}e^{-H[\{\sigma_{i}\}]/2})^{2}\geq\sum_{\{\sigma_{i}\}}e^{-H[\{\sigma_{i}\}]}$, we see that the right hand side of (71) is at least as big as $\sqrt{d^{\prime}\mathcal{Z}_{+-}}$, as claimed in the main text.
@@
Note that the when calculating quasientropy, the expression one obtains [Eq. (21)] is a ratio of partition functions with different boundary conditions $\mathcal{Z}_{-+}/\mathcal{Z}_{++}$. There, the bulk contribution to the free energy cancels between the numerator and denominator. In contrast, the bounds we obtain here feature no such denominator. The two agree in the limit where $S_{e}^{(2)}$ scales with system size, because in this case $\mathcal{Z}_{++}$ is dominated by the confiuration where $\sigma_{i}=+1$ for all $i$, which gives $\mathcal{Z}_{++}\approx 1$.
@@ Appendix C Contractivity of measurement channels
In this appendix, we prove contractivity properties of the maps $\tilde{\Theta}_{i}$ [Eq. (59)] for generalized measurements. We focus on discrete POVMs for simplicity, but the following generalizes to continuous POVMs straightforwardly. Let $\{F_{s}\}$ be the collection of POVM operators that define the measurement process $\mathcal{T}_{i}$, which satisfy $\sum_{s}F_{s}=\mathbbm{I}$, and let $\sigma_{i}$ be the state appearing in definition (59), which we are free to choose. For an arbitrary operator $X$ on $Q_{i}$, we want to find an upper bound for
@@
$\min_{\sigma_{i}}\|\tilde{\Theta}_{i}[X]\|_{2}^{2}=\min_{\sigma_{i}}\operatorname{Tr}\Big{[}\sigma_{i}^{-1/2}\Theta_{i}[X]\sigma_{i}^{-1/2}\Theta_{i}[X]\Big{]}$ (73)
in terms of the norm $\|X\|_{2}$. Since the output of $\Theta_{i}$ is classical (i.e. diagonal), any choice of $\sigma_{i}$ that achieves this minimum will also be diagonal [117], thus we can set $\sigma_{i}=\text{diag}(\{q_{s}\})$ for some set of $q_{s}\geq 0$ satisfying $\sum_{s}q_{s}=1$. This gives
$\min_{\sigma_{i}}\|\tilde{\Theta}_{i}[X]\|_{2}^{2}=\min_{q_{s}}\sum_{s}q_{s}^{-1}\operatorname{Tr}[F_{s}\cdot(X-\operatorname{Tr}[X]\pi^{i})]^{2}$ (74)
In principle, we could consider choosing $q_{s}$ in a way that depends on $X$, Because the minimum on the right hand side can be upper bounded by taking any fixed choice for $q_{s}$, from this point on we will take
$q_{s}=\operatorname{Tr}[F_{s}]/d_{Q_{i}}$ (75)
where $d_{Q_{i}}$ is the Hilbert space dimension of $Q_{i}$ (which is equal to $q$ for most of this work). This choice is motivated by the interpretation of POVMs as ensembles of states wherein the density matrix $\omega_{s}\coloneqq F_{s}/\operatorname{Tr}[F_{s}]$ occurs with probability $q_{s}$ [95]. We will refer to this ensemble as the POVM ensemble in the following. This specific choice gives us an upper bound on the true minimum in (73), which is what we desire.
[…]
where $\text{where }\hat{\Phi}\coloneqq$ $\;\mathcal{P}_{0}\circ\Phi\circ\mathcal{P}_{0},$
[…]
where $\mathcal{P}_{i}=\text{id}_{i}-\mathcal{D}^{Q_{i}}$ is the projector onto the space of traceless operators on site $i$. (The map $\Phi$ arises naturally in the theory of state tomography [118, 119, 120].) Evidently, we can claim that the map $\tilde{\Theta}_{i}$ for our chosen value of $q_{s}$ (75) obeys
$\|\tilde{\Theta}_{i}[X]\|_{2}$ $\leq\lambda_{i}\|X\|_{2}$ $\forall X$ (78)
with contractivity constant
$\lambda_{i}\leq\|\hat{\Phi}\|_{\infty}^{1/2}.$ (79)
Note that the map $\Phi$ encodes all the second moments of the POVM ensemble. Hence, for 2-design measurements, the map $\Phi$ can be computed using a Haar integral, which yields $\Phi=(\text{id}+\mathcal{D}^{Q_{i}})/[d_{i}(d_{i}+1)]$, and in turn we have $\lambda_{i}\leq 1$. This proves Eq. (61), including in cases where the POVM cannot be written as a unitary 2-design followed by a projective measurement.
@@
Note that due to the prefactor of $d_{i}^{2}$ in Eq. (77), if we generalize to measurement operations for which the POVM ensemble forms a $\epsilon$-approximate state design, we can estimate the size of the correction correction to the contractivity constant $\lambda_{i}$ as being $\epsilon d_{i}^{2}$. This may be somewhat larger than one might naïvely expect in cases where $\epsilon$ is taken to have a constant magnitude—an issue that was noted in the context of decoupling problems in Ref. [121]. Unfortunately, this means we cannot directly apply powerful results on the formation of designs in shallow quantum circuits [122, 123], which is why we require a more specialised analysis the spatial structure of entanglement in these circuits (Appendix E).
@@ Appendix E Detailed bound for random 4-local circuits
Evidently, the POVM describing the effective measurement process within each cell does not form a 2-design, but rather a tensor product of eight separate 2-designs, corresponding to each of the eight four-site gates applied immediately before measurement. Our starting point will be Eq. (55), which we obtained by considering the coarse-grained cells, but now we will ‘fine-grain’ again and separate out each group of four qudits that share the same gate in the final layer. We label each quartet within a cell $i$ by an index $a=1,\ldots,8$. We can use the same approach as Eq. (39) to split $\Theta_{i}$ into a sum of $2552^{8}-1=$ 255 terms

### v2

@@ VII.2 Comparison to measurement-based quantum computation
The cluster state—the canonical resource for performing measurement-based quantum computation (MBQC) [7]—is a specific example of a state that can be prepared in constant depth and for which we can guarantee long-ranged MIE (provided we measure in the appropriate basis): By construction, this state has the property that upon measuring qubits in bases within the $x$-$y$ plane, a random unitary circuit is executed in the sideways direction. The specific unitary implemented in this sideways circuit depends on the measurement outcomes $s_{B}$, and thus the post-measurement state received by $A$ and $C$ after this process is given by |$\ket{\phi^{AC}_{s_{B}}}=(I_{A}\otimes U^{\rm MBQC}_{s_{B}})\ket{\Phi^{+}_{AC}}$, where $U^{\rm MBQC}_{s_{B}}$ describes the unitary circuit conditioned on the outcomes, and |$\ket{\Phi^{+}_{AC}}$ is some maximally entangled state. We can relate the unitarity of this sideways circuit to the fact that such states have maximal measurement-induced entanglement (this is a consequence of the Schrödinger-HJW theorem [95]).
@@
We remark that the picture described here very naturally applies to the problem of measuring cluster states in bases out of the $x$-$y$ plane, where entanglement transitions have been observed numerically [16, 81]. Additionally, the sideways-evolving circuit picture provides a new way to analyse strategies for distributing entanglement in quantum networks [103, 104], where similar kinds of short- to long-ranged entanglement transitions can occur.
@@ VII.3 Connections to hardness of sampling
As mentioned in the introduction, there is a close relationship between measurement-induced entanglement in a given setting and the hardness of simulating the output distribution of the underlying state [15, 19, 105]. Indeed, the known worst-case examples of finite-depth states that are provably hard to sample from [2, 106] (assuming certain widely-believed conjectures in complexity theory) use resource states for MBQC, which as we saw above can exhibit maximal MIE. Here we discuss how our results can be placed within the context of quantum advantage in shallow-depth circuits.
@@ VIII Conclusion and outlook
Our results raise several important open questions. One natural direction would be to consider the effects of experimentally relevant perturbations, such as noise and measurement readout errors. While a constant rate of noise without any change of circuit architecture is likely to lead to result in a distribution that is easy to sample from classically (as is the case in deep random circuit sampling [112]), hardness results for noisy circuit sampling in the constant-depth regime have been put forward in three-dimensional systems [113, 114], and also in certain adaptive settings [115]. One can also ask the separate question of whether post-measurement states in these settings with noise remain long-ranged entangled, which could be addressed by generalizing our distillation-based arguments to mixed states, e.g. using a decoupling approach [64]. In addition, it would be interesting to consider rigorous approaches to upper bounding MIE, in order to prove the stability of the conjectured low-MIE, classically simulable phase [15].
The states we have considered in this work are by definition short-ranged entangled. Using our isoTNS representation of these states, we showed how the sampling problem is equivalent to a non-unitary monitored quantum dynamics evolving in the spatial direction. More generally, isometric tensor networks can be used to represent a more general class of wavefunctions, including those with topological order [45]. It would be interesting to understand how the structure of such long-ranged entangled states imprints itself on the effective monitored dynamics, and in turn, how the post-measurement states are themselves affected, particularly given recent results linking topological phases with particular patterns of MIE [116].
More generally, there are a wide range of different settings in which such transitions in MIE can occur, and a more detailed understanding of the physics of these phenomena would shed new light on the role of entanglement in many-body quantum dynamics. There is scope for more extensive numerical studies of these transitions, e.g. to better pinpoint the location and nature of the critical point in different circuit architectures, which would also help quantify the tightness of our bounds. Similarly, we remark on the possibility of experimental studies of these phenomena, building on recent realisations of shallow circuit sampling in trapped ion quantum computers [117]. While detecting and verifying MIE can be challenging due to the randomness of the post-measurement state, a number of recent strategies have been put forward which could be used to directly observe these transitions [86, 87].
@@ Acknowledgements.
MM acknowledges support from Trinity College, Cambridge. WWH is supported by the Singapore NRF Fellowship, NRF-NRFF15-2023-0008, and through the National Quantum Office, hosted in A*STAR, under its Centre for Quantum Technologies Funding Initiative (S24Q2d0009). DM acknowledges support from the Novo Nordisk Fonden under grant number NNF22OC0071934 and NNF20OC0059939. Discussions between MM and WWH on the subject of this work were initiated in part at the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-1607611.
@@ Appendix A Isometric tensor network representation of finite-depth states
Figure 10: (a) After grouping legs together appropriately, the collections of gates depicted in Fig. 9 can be connected up into the network shown. Black dots representing physical (blocked) qudits, which are arranged into cells, demarcated with red dashed lines. The intra-cell unitaries are left implicit here. (b) If we consider all qudits within a cell as one collective degree of freedom, we can write |$\ket{\Psi_{\rm 2D}}$ as a tensor network state, where each five-legged tensor (c, blue squares) is an isometry from two virtual legs to the physical leg (out of the page) plus the other two virtual legs. This defines an isoTNS state [46], with an orthogonality center in a particular corner of the lattice.
@@ Appendix B Ising partition function bound for holographic model
where [$\mathcal{H}[\{\sigma_{i}\}]$ is equal to the 2-Rényi entropy of the corresponding region $I$ in the pre-measurement state, as we saw in Eq. (22). In the holographic circuit model, $H[\{\sigma_{i}\}]$ is equal to the entropy of the bond state $S_{e}^{(2)}$ times the total number of bonds on which the Ising spins anti-align. The right hand side of (73) then becomes the partition function for the Ising model with opposite boundary conditions on $A$ and $C$, where the inverse temperature is $\beta=S_{e}^{(2)}/2$. We can compare this to $\mathcal{Z}_{+-}$ [Eq. (21)], which we encountered in the calculation of quasientropy: the difference is just a factor of 2 in the effective temperature. By the Cauchy-Schwartz inequality ($(\sum_{\{\sigma_{i}\}}e^{-H[\{\sigma_{i}\}]/2})^{2}\geq\sum_{\{\sigma_{i}\}}e^{-H[\{\sigma_{i}\}]}$, we see that the right hand side of (73) is at least as big as $\sqrt{d^{\prime}\mathcal{Z}_{+-}}$, as claimed in the main text.
@@
Note that when calculating quasientropy, the expression one obtains [Eq. (21)] is a ratio of partition functions with different boundary conditions $\mathcal{Z}_{-+}/\mathcal{Z}_{++}$. There, the bulk contribution to the free energy cancels between the numerator and denominator. In contrast, the bounds we obtain here feature no such denominator. The two agree in the limit where $S_{e}^{(2)}$ scales with system size, because in this case $\mathcal{Z}_{++}$ is dominated by the confiuration where $\sigma_{i}=+1$ for all $i$, which gives $\mathcal{Z}_{++}\approx 1$.
@@ Appendix C Contractivity of measurement channels
In this appendix, we prove contractivity properties of the maps $\tilde{\Theta}_{i}$ [Eq. (55)] and $\tilde{\Theta}_{A}$ [Eq. (58)], following the derivations given in Refs. [62, 27]. For convenience, we recall the definition of the distillation channel applied to $A$
$\mathcal{T}^{A\rightarrow A^{\prime}M_{A}}[\rho^{A}]=\frac{d_{A}}{d^{\prime}}\sum_{r_{A}}q_{r_{A}}L_{r_{A}}^{\dagger}\rho^{A}L_{r_{A}}\otimes\ket{{r_{A}}}\!\!\bra{{r_{A}}}_{M_{A}},$ (75)
in terms of which $\Theta_{A}\coloneqq(\text{id}^{A^{\prime}}-\mathcal{D}^{A^{\prime}})\circ\mathcal{T}^{A\rightarrow A^{\prime}M_{A}}$. Here, $L_{r_{A}}^{\dagger}=L_{0}^{\dagger}V_{r_{A}}$, where $L_{0}$ is a fixed isometry from $A^{\prime}$ to $A$, and $\{(q_{r_{A}},V_{r_{A}})\}$ is a unitary 2-design. Because the measurement channels on $B$ are a special case of the above channel for $d^{\prime}=1$, we will prove the more general bound (58) here.
For an arbitrary operator $X$ on $A$, we want to find an upper bound for
$\min_{\sigma^{A^{\prime}M_{A}}}\|\tilde{\Theta}_{A}[X]\|_{2}^{2}$
$=$ $\min_{\sigma^{A^{\prime}M_{A}}}\operatorname{Tr}\left[(\sigma^{A^{\prime}M_{A}})^{-1/2}\Theta_{i}[X](\sigma^{A^{\prime}M_{A}})^{-1/2}\Theta_{i}[X]\right]$ (76)
where the minimum is over all density matrices on $A^{\prime}M_{A}$. Finding this minimum is in general challenging; however, we can get the required upper bound by using any fixed choice of $\sigma_{A^{\prime}M_{A}}$. In particular, we choose
$\sigma^{A^{\prime}M_{A}}=\pi^{A^{\prime}}\otimes\text{diag}(q_{r})^{M_{A}}.$ (77)
We then obtain
@@
$\min_{\sigma^{A^{\prime}M_{A}}}\|\tilde{\Theta}_{A}[X]\|_{2}^{2}$ $\leq\frac{d_{A}^{2}}{d^{\prime}}\sum_{r_{A}}q_{r_{A}}\bigg{(}\operatorname{Tr}[(L_{0}^{\dagger}V_{r_{A}}XV_{r_{A}}L_{0})^{2}]$
$-\frac{1}{d^{\prime}}\operatorname{Tr}[L_{0}^{\dagger}V_{r_{A}}XV_{r_{A}}L_{0}]^{2}\bigg{)},$ (78)
where $\Pi_{0}\coloneqq L_{0}L_{0}^{\dagger}$ is a projector of rank $d^{\prime}$. Noticing that the above involves only second moments of the ensemble of unitaries $\{(q_{r_{A}},V_{r_{A}})\}$, we can us the 2-design property to replace the average with an integral with respect to the Haar measure ${\rm d}\mu_{H}(U_{A})$ over $U_{A}\in\mathrm{U}(d_{A})$
$\min_{\sigma^{A^{\prime}M_{A}}}\|\tilde{\Theta}_{A}[X]\|_{2}^{2}$
$\leq\frac{d_{A}^{2}}{d^{\prime}}\int{\rm d}\mu_{H}(U_{A})\operatorname{Tr}[(\mathcal{F}_{A^{\prime}}-\mathcal{I}_{A^{\prime}}/d^{\prime})\cdot(L_{0}^{\dagger}U_{A}XU_{A}^{\dagger}L_{0})^{\otimes 2}],$ (79)
where we have used the swap operator $\mathcal{F}_{A^{\prime}}$ to express the first term using the identity $\operatorname{Tr}[(Y_{A^{\prime}})^{2}]=\operatorname{Tr}[\mathcal{F}_{A^{\prime}}Y_{A^{\prime}}^{\otimes 2}]$, and $\mathcal{I}_{A^{\prime}}\coloneqq I_{A^{\prime}}\otimes I_{A^{\prime}}$. Now all that remains is to evaluate this integral with the help of Weingarten calculus [120]. Noting that $\operatorname{Tr}[(\mathcal{F}_{A^{\prime}}-\mathcal{I}_{A^{\prime}}/d^{\prime})\cdot(L_{0}L_{0}^{\dagger})^{\otimes 2}]=0$, we can use the following expression which holds for traceless $Y$
$\int{\rm d}\mu_{H}(U_{A})\operatorname{Tr}[YU^{\otimes 2}_{A}X(U^{\dagger}_{A})^{\otimes 2}]$
$=$ $\frac{1}{d_{A}^{2}-1}\Bigg{[}\operatorname{Tr}[Y^{2}]\operatorname{Tr}\Big{[}\big{(}(\text{id}^{A}-\mathcal{D}^{A})[X]\big{)}^{2}\Big{]}\Bigg{]}$ if $\text{if }\operatorname{Tr}[Y]=0$ (80)
Identifying $Y=L_{0}^{\otimes 2}(\mathcal{F}_{A^{\prime}}-\mathcal{I}_{A^{\prime}}/d^{\prime})(L_{0}^{\dagger})^{\otimes 2}$, we obtain
$\min_{\sigma^{A^{\prime}M_{A}}}\|\tilde{\Theta}_{A}[X]\|_{2}^{2}$ $\leq\frac{d_{A}^{2}}{d^{\prime}}\frac{(d^{\prime})^{2}-d^{\prime}/d_{A}}{d_{A}^{2}-1}\|(\text{id}^{A}-\mathcal{D}^{A})[X]\|_{2}^{2}$
$\leq d^{\prime}\|X\|_{2}^{2},$ (81)
which gives us Eq. (58). Since the measurement channel for each qudit in $B$ is equivalent to the distillation channel (75) for the case of fully destructive measurements $d^{\prime}=1$, we immediately obtain Eq. (57) as a corollary.
@@
Note that due to the prefactor of $d_{A}^{2}/d^{\prime}$ in Eq. (79), if we generalize to measurement operations for which the POVM ensemble forms a $\epsilon$-approximate state design, we can estimate the size of the correction correction to the contractivity constant $\lambda_{i}$ as being $\epsilon d_{A}^{2}/d^{\prime}\geq\epsilon d_{A}$. This may be somewhat larger than one might naïvely expect in cases where $\epsilon$ is taken to have a constant magnitude—an issue that was noted in the context of decoupling problems in Ref. [121]. Unfortunately, this means we cannot directly apply powerful results on the formation of approximate designs in shallow quantum circuits [122, 123], which is why we require a more specialised analysis the spatial structure of entanglement in these circuits (Appendix E).
@@ Appendix D Summing over domain walls
As a technical note, if the original lattice $\Lambda$ is a planar graph that is not a triangulation, then we have the added complication that there can be faces where domain walls intersect. As explained previously, we can always add edges to $\Lambda$ to make a triangulation $\Lambda^{\prime}$, while ensuring that the strictly local condition Definition 4 still holds with respect to $\Lambda^{\prime}$. This will yield a bound on $\overline{\epsilon}$ in terms of self-avoiding walks on the dual lattice of $(\Lambda^{\prime})^{*}$, which may be distinct from $\Lambda^{*}$. For each self-avoiding walk on $(\Lambda^{\prime})^{*}$, there is a corresponding self-avoiding walk on $\Lambda^{*}$, and because the Boltzmann weights only depend on the Rényi entanglement entropy across $W$ (which does not change when extra fictitious bonds are added), we can substitute the partition function for the original dual lattice $\Lambda^{*}$.
@@ Appendix E Detailed bound for random 4-local circuits
Evidently, the POVM describing the effective measurement process within each cell does not form a 2-design, but rather a tensor product of eight separate 2-designs, corresponding to each of the eight four-site gates applied immediately before measurement. While Eq. (51) was obtained by considering coarse-grained cells, we will now have to ‘fine-grain’ again and separate out each quartet of four qudits that share the same gate in the final layer, of which there are 8 per coarse-grained cell. We label each quartet within a cell $i$ by an index $a=1,\ldots,8$, so we can write $\mathcal{T}^{Q_{i}\rightarrow M_{i}}=\bigotimes_{a=1}^{8}\mathcal{T}_{i,a}$, where $\mathcal{T}_{i,a}$ are the measurements on a single quartet $a$. We can use the same approach as Eq. (39) to split $\Theta_{i}$ into a sum of $2552^{8}-1=$ 255 terms